Find the values of x, y, and z that maximize xyz subject to the constraint 192-x-16y – 2z = 0.
Answer #1

flogy,z)=ky Z f = 79. Let Giwer constraint Constraint gx, 192-x-164-22=0 Let g(ny, z) = 192-1-164-22 By lagrange lagrange mebof=<yz, xz, xy) 79 = _923 Sy, 92 > - < (192 - x - 164 -22) (192 - x - 164-22), 22 (192-x-164-223) زہ a dy (-, -16, -27 ,. Tg= 2z=164 z = sy x=2Z, y = 2 Consider giver constraint 192 - x 16y. 2z=0 192- 22 -1673)-27 = -) - 2Z - 22 - 22=492 = z = 192 6